A thick walled hollow sphere has outer radius R. It rolls down an inclined plane without slipping and its speed at the bottom is u. If the inclined plane is frictionless and the sphere s!ides down without rolling, its speed at the bottom will be 5u/ 4 . What is the radius of gyration of the sphere?

Grade:12

1 Answers

J.s. Bharatraj
6 years ago
Initially, when the sphere is rolling down the include without any slipping, then let the angular velocity of the sphere at the bottom of the incline be w. 
 
No slipping condition, V = wR
Furthermore, due to conservation of energy
PE lost = KE gained
mgh = 1/2mv^2 + 1/2 *I*w^2 (Where h is the height of the incline and I is the moment of inertia of sphere around its center)
Also, I = m*Rg ^2  where Rg is the radius of gyration and also v = wR (No slipping condition)
Hence, mgh = 1/2mv^2 + 1/2 *m*Rg ^2 *v^2 /R^2
2gh = [1+(Rg/R)^2]*v^2 (1)
 
Furthermore, under no rolling case
mgh = 1/2m*(5v/4)^2
v^2 = 32/25 gh
Putting it in the equation 1
2gh = [1+(Rg/R)^2]*32/25 gh
Hence, Rg = 3/4R
So, the radius of gyration of the sphere around its center is 3/4R